OpenStudy (diyadiya):

A 5kg stone falls from a height of 100m and penetrates 2m in a layer of sand. The time of penetration is ?

OpenStudy (diyadiya):

answer is 0.29sec

OpenStudy (anonymous):

ok...one min..

OpenStudy (diyadiya):

sure

OpenStudy (anonymous):

is it dropped cheche?

OpenStudy (diyadiya):

It falls from a height ,thats what given..

OpenStudy (anonymous):

velocity during approach to sand is 44.72m/s (use formula V=sqrt(2gh) Vi=44.72,Vf=0 v=u-at t=(u-v)/a=44.72/a s=0.5*a*t^2=0.5*(44.72/t)*t^2 4=44.72t t=4/44.72 = 0.089s,is it to be done like this????

OpenStudy (anonymous):

ya tahts exactly wat i got

OpenStudy (anonymous):

so i dont get whats wrong???

OpenStudy (apoorvk):

ahaa!! now see. the stone has achieved a max velocity (2gh)^(1/2) just before it hits the ground. and decelerates to 0, while it travels 2m in the sand so you have initial and final velocity, distance travelled, find deceleration rate. then find the time taken for that decelration!!

OpenStudy (anonymous):

rama that wat we did!

OpenStudy (apoorvk):

LOLLOL

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

taht mass=5kg data gives us no substantial use

OpenStudy (mos1635):

s=0.5*a*t^2=0.5*(44.72/t)*t^2 Sarkar that is not so S=Vo*t - 1/2 *a* t^2 t=Vo/a and so S=Vo^2/2a

OpenStudy (anonymous):

yes i thought of that@salinl

OpenStudy (apoorvk):